OpenStudy (jhannybean):

Calculus 3: Description of inequalities of the solid above \(z=\sqrt{x^2+y^2}\) and below \(x^2+y^2+z^2=z\)

OpenStudy (jhannybean):

I understand that because the solid is ABOVE the cone, \(z\ge \sqrt{x^2+y^2} \implies z^2 \ge x^2+y^2\)

OpenStudy (jhannybean):

@.Sam. or @ganeshie8 ? :)

OpenStudy (jhannybean):

How do you get from what I have in my first post to \(2z^2 \ge x^2+y^2+z^2 = \rho^2\) ?

OpenStudy (andrewthecookie):

Symbols hurting my brain >_<

OpenStudy (jhannybean):

Sorry :(

OpenStudy (andrewthecookie):

Its k :P

OpenStudy (greencat):

I'm still in College Algebra. What does this have to do with solid?

OpenStudy (jhannybean):

You're trying to find a volume of a solid by understanding it's coordinates on a 3D axis :)

OpenStudy (greencat):

x^2 + y^2 + z^2 = z ===> x^2 +y^2 = 1/z

OpenStudy (greencat):

can this be solved algebraicly?

OpenStudy (jhannybean):

Possibly, but there's a substitution involved and I don't know how to get there :(

OpenStudy (greencat):

it seems as i z^2 = 1/z

OpenStudy (greencat):

if*

OpenStudy (greencat):

becuz of the x^2 +y^2 is in both equations

OpenStudy (greencat):

That is if x y and z are variables. if the x y nd z of both equations are different then im wrong

OpenStudy (greencat):

Im sorry fot speaking without knowing anything.

OpenStudy (greencat):

for*

OpenStudy (greencat):

Cat be sorry.

zepdrix (zepdrix):

Grr Calc 3 was so hard :(

zepdrix (zepdrix):

If you fiddle with the lower boundary a little bit,\[\Large\rm x^2+y^2+\left(z-\frac{1}{2}\right)^2=\left(\frac{1}{2}\right)^2\]You can see that we have a sphere of radius 1/2 which has been shifted up 1/2. I don't know if that helps us though >.< So it looks like they're converting to Spherical Coodinates maybe?

zepdrix (zepdrix):

blah, hmm

OpenStudy (jhannybean):

Sorry, was afk trying to figure this out.

OpenStudy (jhannybean):

@zepdrix , I have it partially worked out, can you explain some of my steps? Haha.

OpenStudy (jhannybean):

Ok from the start.

OpenStudy (jhannybean):



OpenStudy (jhannybean):

it lies above the cone.

OpenStudy (jhannybean):

above cone, \( z= \sqrt{x^2+y^2}\) and below he sphere \(x^2 +y^2 +z^2 = z\) That means...\[z \ge \sqrt{x^2+y^2} \implies z^2\ge x^2+y^2\]

OpenStudy (tkhunny):

WHAT are you doing? Have you considered Polar Coordinates?